# thanks: Corresponding author

### v2

@@ Abstract
Introduction.As a gravitational system with thermodynamic properties, black holes exhibit rich phase structures. In particular, phase transitions in anti-de Sitter (AdS) spacetime offer profound insights into the nature of quantum gravity. Significant examples include the Hawking-Page phase transition between thermal AdS and large black hole phases Hawking:1982dh and the van der Waals-type phase transition in charged AdS black holes Kubiznak:2012. These discoveries reveal that black holes possess thermodynamic behavior as rich as that of ordinary matter and are intimately connected to key phenomena in condensed-matter physics through the AdS/CFT correspondence.
Recent work has increasingly explored these phase transitions from geometric and topological perspectives, providing new insight into this important field. It has been demonstrated that the synchronized multivalued behavior of dynamical quantities such as the Lyapunov exponent and geometric quantities such as the intrinsic or extrinsic curvature near the phase transition point can diagnose a first-order phase transition Guo:2022; Yang:2023; Lyu:2024; Kumara:2024; Du:2025; Shukla:2024; Gogoi:2024; Chen:2025xqc; R:2025gok; Awal:2025irl; Yang:2025; Kumar:2025kzt; Guo:2025pit; Bezboruah:2025udi; Ali:2025ooh; Zhang:2025; Xie:2025auj; Zhang:2025kqd. Moreover, since the Lyapunov exponent is quantitatively linked to geometric quantities in both null and timelike cases, this connection has established a unified perspective linking dynamics, spacetime geometry, and first-order phase transitions Gallo:2025; Zhang:2025; Zhang:2025kqd. Separately, topological classification schemes have revealed that black holes can be categorized according to their topological properties Wei:2021vdx; Wei:2022dzw; Wei:2024gfz, offering new insights for a deeper understanding of the fundamental nature of quantum gravity.
@@
Based on these considerations, we propose a universal geometric criterion for first-order phase transitions: examining whether the curve of the temperature function $T(r_{+})$ possesses two extrema (a local maximum and a local minimum). Equivalently, this criterion requires that $\partial T/{\partial r_{+}}=0$ admits two distinct positive real roots. This criterion provides an intuitive and computable diagnostic tool that enables a clear classification based on the local geometric structure underlying the phase transition. Our work demonstrates that the spacetime structure governed by the Einstein equations induces specific geometric features, such as non-degenerate critical points in the thermodynamic parameter space of a black hole. This results in quantities such as the photon sphere radius, Lyapunov exponent, and intrinsic/extrinsic curvatures exhibiting multivalued or non-monotonic dependence on temperature $T$ Guo:2022; Yang:2023; Lyu:2024; Kumara:2024; Du:2025; Shukla:2024; Gogoi:2024; Chen:2025xqc; R:2025gok; Awal:2025irl; Yang:2025; Kumar:2025kzt; Guo:2025pit; Bezboruah:2025udi; Ali:2025ooh; Zhang:2025; Xie:2025auj; Zhang:2025kqd; PS01; PS02; PS03; PS04; PS05; PS06; PS07; PS08; PS09. These features intrinsically encode the information of phase transitions. Therefore, the geometry of the thermodynamic parameter space offers a profound and unified perspective for understanding black hole phase transitions. We set $G=c=k_{B}=\hbar=1$ in this Letter.
@@
Considering an unstable null circular orbit (light ring) at radius $r_{LR}$ near the black hole, its Gaussian curvature is given by Gallo:2025
@@
The multivalued behavior of these quantities, which reflect chaotic dynamics and spacetime geometry, shows a precise consistency with the thermodynamic free energy $\tilde{F}$. This indicates that phase transition information is universally encoded in multiple properties of spacetime Zhang:2025; Zhang:2025kqd. In the following analysis, we will investigate this characteristic phenomenon and uncover the deep mathematical structure and physical mechanisms behind it.
@@
This conclusion has a deeper geometric counterpart in the theory of covering spaces. At each critical point $r_{v}$ (with $v=1,2$), the Morse lemma guarantees that $T(r_{+})$ can be locally expressed as $T\approx T_{v}-y^{2}$ for a local maximum and $T\approx T_{v}+y^{2}$ for a local minimum, where $y$ measures the deviation from the non-degenerate critical point Morse:1925. These points are therefore fold singularities, which directly implies that near $T_{1}$ or $T_{2}$, a single temperature corresponds to two distinct horizon radii $r_{+}$ (one double root and one distinct root). This folding of the parameter space is formalized by introducing a covering manifold. We define a one-dimensional manifold $\mathcal{M}$ parameterized by all horizon radii $r_{+}\in I\subset\mathbb{R}$ that correspond to such black hole solutions. We accordingly construct a three-sheeted covering manifold $\widetilde{\mathcal{M}}=J_{1}\sqcup J_{2}\sqcup J_{3}$, where $J_{1}=(\mathcal{A},\,r_{1}),~J_{2}=(r_{1},\,r_{2}),~J_{3}=(r_{2},\,\mathcal{B})$ with $\mathcal{A}<r_{1}<r_{2}<\mathcal{B}$. Here, $\mathcal{A}$ and $\mathcal{B}$ are points in the domain $I\subset\mathbb{R}^{+}$, and the interval $(\mathcal{A},\,\mathcal{B})$ constitutes the relevant parameter range for the phase transition. The projection $\widetilde{\pi}$$\widetilde{\mathcal{M}}\rightarrow\mathbb{R}$, defined by $\widetilde{\pi}(r_{+})=T(r_{+})$, is injective on each branch. The topological structure of the base manifold $\mathcal{M}$ ensures that its covering space $\widetilde{\mathcal{M}}$ must be three-sheeted on the spinodal region. This explains the geometric origin of multivaluedness within the black hole parameter manifold.
@@
Consequently, any physical quantity $\mathcal{F}$ that depends on $r_{+}$ inherits this multivaluedness. Explicitly, $\mathcal{F}(T)$ can be regarded as the composition of its lift $\mathcal{F}^{\uparrow}$ on the covering space with the inverse projection $\widetilde{\pi}^{-1}(T)$. Since $\widetilde{\pi}^{-1}(T)$ yields three distinct $r_{+}$ values and $\mathcal{F}^{\uparrow}$ is not identically equal on different branches, $\mathcal{F}(T)$ naturally exhibits multivaluedness. Conversely, observing multivaluedness in $\mathcal{F}(T)$ implies the existence of non-degenerate critical points in $T(r_{+})$, thereby diagnosing a first‑order phase transition. This explains why the use of multivaluedness to detect first‑order phase transitions in Refs. Guo:2022; Yang:2023; Lyu:2024; Kumara:2024; Du:2025; Shukla:2024; Gogoi:2024; Chen:2025xqc; R:2025gok; Awal:2025irl; Yang:2025; Kumar:2025kzt; Guo:2025pit; Bezboruah:2025udi; Ali:2025ooh; Zhang:2025; Xie:2025auj; Zhang:2025kqd is justified, thereby providing a rigorous mathematical foundation for this method. Complete technical details are given in the Supplemental Materials.
@@
A local geometric criterion for classifying phase transitions.Since $\mathcal{F}(T)$ inherits the multivaluedness of $r_{+}(T)$ during a first-order phase transition, we propose a classification scheme for black holes based on the local geometric criterion introduced above. It categorizes black holes into classes $A1$, $A2$, and $B$ according to the number of local extrema in the $T(r_{+})$ curve. The $A1$ class corresponds to curves with two distinct extrema, signaling two non-degenerate critical points. This leads to a three-valued $\mathcal{F}(T)$, resulting in a first-order phase transition. In contrast, the $A2$ class exhibits only one critical point and therefore does not display the characteristic features of a first-order phase transition. Finally, for the $B$ class, the curve has no local extrema, thus $\mathcal{F}(T)$ is strictly monotonic, which means the absence of a phase transition for such black holes.
@@
As shown in Fig. 2, for an RN–AdS black hole with $Q=(\ell/8.66)<Q_{c}$, the $T(r_{+})$ curve displays a local maximum $T_{1}$ and a local minimum $T_{2}$, signaling a typical first-order phase transition. In contrast, for a series of charges $Q>Q_{c}$ (such as $\ell/5.9$, $\ell/5$, $\ell/4.5$, $\ell/4$), the curves are monotonic with no local extrema, indicating the absence of a first-order phase transition. This result agrees with the conclusion obtained from the thermodynamic analysis Guo:2022, thus providing support for our geometric framework.
@@
The discriminant of this quadratic equation is $\Delta=1-\frac{36Q^{2}}{\ell^{2}}>0$. When $\Delta>0$ (equivalently $\ell>6Q$), the equation has two distinct positive real roots, indicating a first-order phase transition and placing the black hole in the $A1$ class. The condition $\ell>6Q$ matches the critical charge from the standard thermodynamic criticality condition Eq. (4), which confirms the effectiveness of our multivaluedness‑based diagnostic.
@@
The only physical result is $r_{+}=\sqrt{\ell^{2}/{3}}$. Hence, the $T(r_{+})$ curve has only one critical point. The SAdS black hole therefore does not exhibit the characteristics of a first-order phase transition and belongs to the $A2$ class.
@@
Our classification scheme, rooted in the analysis of local geometry in the thermodynamic parameter space, offers a perspective complementary to the framework in Ref. Wei:2024gfz, which is based on global topological invariants. Their work assigns the RN–AdS and SAdS black holes to the topological classes $W^{0+}$ and $W^{0-}$. These two distinct perspectives highlight the depth and effectiveness of the geometric approach in studying black hole thermodynamics.
Extensions and universality of the geometric framework.The geometric mechanism proposed in this Letter is not restricted to spherically symmetric black holes. For rotating black holes such as Kerr‑AdS and Kerr‑Newman‑AdS, the temperature curve $T(r_{+})$ at fixed angular momentum (and charge) also exhibits two local extrema when a first‑order phase transition occurs Chen:2025xqc; Yang:2025. According to our criterion, these rotating black holes therefore belong to the $A1$ class, and the multivaluedness of related dynamical or geometric quantities arises from the same folded geometry.
Furthermore, in Refs. PS01; PS02; PS03; PS04; PS05; PS06; PS07; PS08; PS09, diagnosing first-order phase transitions by observing non-monotonic (multivalued) behavior in the diagram of reduced temperature $T/T_{c}$ versus reduced photon sphere radius $r_{ps}/r_{psc}$ (where $T_{c}$ is the critical temperature and $r_{psc}$ the critical photon sphere radius) is essentially equivalent to checking whether $T(r_{+})$ possesses two non-degenerate critical points. This equivalence holds because the photon sphere radius $r_{ps}$ is a physical quantity that continuously depends on the horizon radius $r_{+}$, and its values do not coincide in different branches. In most known black hole solutions, this dependence is in fact monotonic, which yields a one-to-one correspondence between the branches. Consequently, by taking $\mathcal{F}=r_{ps}$, the multivaluedness of $T(r_{+})$ directly implies that of $T(r_{ps})$, and scaling these physical quantities does not change this fundamental property.
@@
As a direct application of this framework, we propose the following geometric criterion for diagnosing a phase transition: a black hole undergoes a first-order phase transition if and only if the temperature curve $T(r_{+})$ exhibits two local extrema. This criterion is equivalent to testing whether $\partial T/\partial r_{+}=0$ has two distinct positive real roots, providing a universal and intuitive tool. Furthermore, we propose a classification scheme that classifies black holes as $A1$, $A2$, and $B$ according to the number of non-degenerate critical points. This classification, based on local geometric properties, forms a complementary perspective to the scheme presented in Ref. Wei:2024gfz, which relies on global topological invariants.
@@ Appendix A Physical nature and mathematical conditions
The mathematical conditions we introduce are grounded in basic physical facts of a first-order phase transition. First, the temperature as a function $T(r_{+})$ is sufficiently smooth, i.e., $T\in C^{p}(I)$ with $p\geq 2$ and $I\subset\mathbb{R}$. More importantly, during a first-order phase transition, this function exhibits two non-degenerate critical points, namely $T^{\prime}(r_{v})=0$ and $T^{\prime\prime}(r_{v})\neq 0$ for $v=1,2$, which correspond to the two endpoints $T_{1}$ and $T_{2}$ of the spinodal region. Outside the spinodal region, $T(r_{+})$ is strictly monotonic. Finally, $T(r_{+})$ is not a constant function on its entire domain. For a black hole that satisfies the Einstein equations, its Hawking temperature $T(r_{+})$ may inherently satisfy or violate these conditions, which directly determines its phase transition behavior.

### v3

@@ Abstract
Introduction.As a gravitational system with thermodynamic properties, black holes exhibit rich phase structures. In particular, phase transitions in anti-de Sitter (AdS) spacetime offer profound insights into the nature of quantum gravity. Significant examples include the Hawking-Page phase transition between thermal AdS and large black hole phases Hawking and Page (1983) and the van der Waals-type phase transition in charged AdS black holes Kubiznak and Mann (2012). These discoveries reveal that black holes possess thermodynamic behavior as rich as that of ordinary matter and are intimately connected to key phenomena in condensed-matter physics through the AdS/CFT correspondence.
Recent work has increasingly explored these phase transitions from geometric and topological perspectives, providing new insight into this important field. It has been demonstrated that the synchronized multivalued behavior of dynamical quantities such as the Lyapunov exponent and geometric quantities such as the intrinsic or extrinsic curvature near the phase transition point can diagnose a first-order phase transition Guo et al. (2022); Yang et al. (2023); Lyu et al. (2024); Kumara et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); R. et al. (2025); Awal and Phukon (2025); Yang et al. (2025a); Kumar et al. (2025); Guo et al. (2025); Bezboruah et al. (2025); Ali and Kuang (2025); Zhang et al. (2025a); Xie and Yang (2025); Zhang et al. (2025b). Moreover, since the Lyapunov exponent is quantitatively linked to geometric quantities in both null and timelike cases, this connection has established a unified perspective linking dynamics, spacetime geometry, and first-order phase transitions Gallo and Mädler (2025); Zhang et al. (2025a, b). Separately, topological classification schemes have revealed that black holes can be categorized according to their topological properties Wei and Liu (2022); Wei et al. (2022, 2024), offering new insights for a deeper understanding of the fundamental nature of quantum gravity.
@@
Based on these considerations, we propose a universal geometric criterion for first-order phase transitions: examining whether the curve of the temperature function $T(r_{+})$ possesses two extrema (a local maximum and a local minimum). Equivalently, this criterion requires that $\partial T/{\partial r_{+}}=0$ admits two distinct positive real roots. This criterion provides an intuitive and computable diagnostic tool that enables a clear classification based on the local geometric structure underlying the phase transition. Our work demonstrates that the spacetime structure governed by the Einstein equations induces specific geometric features, such as non-degenerate critical points in the thermodynamic parameter space of a black hole. This results in quantities such as the photon sphere radius, Lyapunov exponent, and intrinsic/extrinsic curvatures exhibiting multivalued or non-monotonic dependence on temperature $T$ Guo et al. (2022); Yang et al. (2023); Lyu et al. (2024); Kumara et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); R. et al. (2025); Awal and Phukon (2025); Yang et al. (2025a); Kumar et al. (2025); Guo et al. (2025); Bezboruah et al. (2025); Ali and Kuang (2025); Zhang et al. (2025a); Xie and Yang (2025); Zhang et al. (2025b); Wei and Liu (2018); Wei et al. (2019); Zhang et al. (2019); Xu et al. (2019); Li et al. (2020); Naveena Kumara et al. (2020); Du et al. (2023); Kumar et al. (2024); Yang et al. (2025b). These features intrinsically encode the information of phase transitions. Therefore, the geometry of the thermodynamic parameter space offers a profound and unified perspective for understanding black hole phase transitions. We set $G=c=k_{B}=\hbar=1$ in this Letter.
@@
Considering an unstable null circular orbit (light ring) at radius $r_{LR}$ near the black hole, its Gaussian curvature is given by Gallo and Mädler (2025)
@@
The multivalued behavior of these quantities, which reflect chaotic dynamics and spacetime geometry, shows a precise consistency with the thermodynamic free energy $\tilde{F}$. This indicates that phase transition information is universally encoded in multiple properties of spacetime Zhang et al. (2025a, b). In the following analysis, we will investigate this characteristic phenomenon and uncover the deep mathematical structure and physical mechanisms behind it.
@@
This conclusion has a deeper geometric counterpart in the theory of covering spaces. At each critical point $r_{v}$ (with $v=1,2$), the Morse lemma guarantees that $T(r_{+})$ can be locally expressed as $T\approx T_{v}-y^{2}$ for a local maximum and $T\approx T_{v}+y^{2}$ for a local minimum, where $y$ measures the deviation from the non-degenerate critical point Morse (1925). These points are therefore fold singularities, which directly implies that near $T_{1}$ or $T_{2}$, a single temperature corresponds to two distinct horizon radii $r_{+}$ (one double root and one distinct root). This folding of the parameter space is formalized by introducing a covering manifold. We define a one-dimensional manifold $\mathcal{M}$ parameterized by all horizon radii $r_{+}\in I\subset\mathbb{R}$ that correspond to such black hole solutions. We accordingly construct a three-sheeted covering manifold $\widetilde{\mathcal{M}}=J_{1}\sqcup J_{2}\sqcup J_{3}$, where $J_{1}=(\mathcal{A},\,r_{1}),~J_{2}=(r_{1},\,r_{2}),~J_{3}=(r_{2},\,\mathcal{B})$ with $\mathcal{A}<r_{1}<r_{2}<\mathcal{B}$. Here, $\mathcal{A}$ and $\mathcal{B}$ are points in the domain $I\subset\mathbb{R}^{+}$, and the interval $(\mathcal{A},\,\mathcal{B})$ constitutes the relevant parameter range for the phase transition. The projection $\widetilde{\pi}$$\widetilde{\mathcal{M}}\rightarrow\mathbb{R}$, defined by $\widetilde{\pi}(r_{+})=T(r_{+})$, is injective on each branch. The topological structure of the base manifold $\mathcal{M}$ ensures that its covering space $\widetilde{\mathcal{M}}$ must be three-sheeted on the spinodal region. This explains the geometric origin of multivaluedness within the black hole parameter manifold.
@@
Consequently, any physical quantity $\mathcal{F}$ that depends on $r_{+}$ inherits this multivaluedness. Explicitly, $\mathcal{F}(T)$ can be regarded as the composition of its lift $\mathcal{F}^{\uparrow}$ on the covering space with the inverse projection $\widetilde{\pi}^{-1}(T)$. Since $\widetilde{\pi}^{-1}(T)$ yields three distinct $r_{+}$ values and $\mathcal{F}^{\uparrow}$ is not identically equal on different branches, $\mathcal{F}(T)$ naturally exhibits multivaluedness. Conversely, observing multivaluedness in $\mathcal{F}(T)$ implies the existence of non-degenerate critical points in $T(r_{+})$, thereby diagnosing a first‑order phase transition. This explains why the use of multivaluedness to detect first‑order phase transitions in Refs. Guo et al. (2022); Yang et al. (2023); Lyu et al. (2024); Kumara et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); R. et al. (2025); Awal and Phukon (2025); Yang et al. (2025a); Kumar et al. (2025); Guo et al. (2025); Bezboruah et al. (2025); Ali and Kuang (2025); Zhang et al. (2025a); Xie and Yang (2025); Zhang et al. (2025b) is justified, thereby providing a rigorous mathematical foundation for this method. Complete technical details are given in the Supplemental Materials.
@@
A local geometric criterion for classifying phase transitions.Since $\mathcal{F}(T)$ inherits the multivaluedness of $r_{+}(T)$ during a first-order phase transition, we propose a classification scheme for black holes based on the local geometric criterion introduced above. It categorizes black holes into classes $A1$, $A2$, and $B$ according to the number of local extrema in the $T(r_{+})$ curve, as shown in Table 1. The $A2$ class corresponds to curves with two distinct extrema, signaling two non-degenerate critical points. This leads to a three-valued $\mathcal{F}(T)$, resulting in a first-order phase transition. In contrast, the $A1$ class exhibits only one critical point and therefore does not display the characteristic features of a first-order phase transition. Finally, for the $B$ class, the curve has no local extrema, thus $\mathcal{F}(T)$ is strictly monotonic, which means the absence of a phase transition for such black holes.
extrema first-order phase transition branch
$A1$ 1 no 2
$A2$ 2 yes 3
$B$ no no 1
Table 1: Black hole classification based on the number of extrema of the $T(r_{+})$ curve.
@@
As shown in Fig. 2, for an RN–AdS black hole with $Q=(\ell/8.66)<Q_{c}$, the $T(r_{+})$ curve displays a local maximum $T_{1}$ and a local minimum $T_{2}$, signaling a typical first-order phase transition. In contrast, for a series of charges $Q>Q_{c}$ (such as $\ell/5.9$, $\ell/5$, $\ell/4.5$, $\ell/4$), the curves are monotonic with no local extrema, indicating the absence of a first-order phase transition. This result agrees with the conclusion obtained from the thermodynamic analysis Guo et al. (2022), thus providing support for our geometric framework.
@@
The discriminant of this quadratic equation is $\Delta=1-\frac{36Q^{2}}{\ell^{2}}>0$. When $\Delta>0$ (equivalently $\ell>6Q$), the equation has two distinct positive real roots, indicating a first-order phase transition and placing the black hole in the $A2$ class. The condition $\ell>6Q$ matches the critical charge from the standard thermodynamic criticality condition Eq. (4), which confirms the effectiveness of our multivaluedness‑based diagnostic.
@@
The only physical result is $r_{+}=\sqrt{\ell^{2}/{3}}$. Hence, the $T(r_{+})$ curve has only one critical point. The SAdS black hole therefore does not exhibit the characteristics of a first-order phase transition and belongs to the $A1$ class.
@@
Our classification scheme, rooted in the analysis of local geometry in the thermodynamic parameter space, offers a perspective complementary to the framework in Ref. Wei et al. (2024), which is based on global topological invariants. Their work assigns the RN–AdS and SAdS black holes to the topological classes $W^{0+}$ and $W^{0-}$. These two distinct perspectives highlight the depth and effectiveness of the geometric approach in studying black hole thermodynamics.
Extensions and universality of the geometric framework. The geometric mechanism proposed in this Letter is not restricted to spherically symmetric black holes. For rotating black holes such as Kerr‑AdS and Kerr‑Newman‑AdS, the temperature curve $T(r_{+})$ at fixed angular momentum (and charge) also exhibits two local extrema when a first‑order phase transition occurs Chen et al. (2025); Yang et al. (2025a). According to our criterion, these rotating black holes therefore belong to the $A2$ class, and the multivaluedness of related dynamical or geometric quantities arises from the same folded geometry.
Furthermore, in Refs. Wei and Liu (2018); Wei et al. (2019); Zhang et al. (2019); Xu et al. (2019); Li et al. (2020); Naveena Kumara et al. (2020); Du et al. (2023); Kumar et al. (2024); Yang et al. (2025b), diagnosing first-order phase transitions by observing non-monotonic (multivalued) behavior in the diagram of reduced temperature $T/T_{c}$ versus reduced photon sphere radius $r_{ps}/r_{psc}$ (where $T_{c}$ is the critical temperature and $r_{psc}$ the critical photon sphere radius) is essentially equivalent to checking whether $T(r_{+})$ possesses two non-degenerate critical points. This equivalence holds because the photon sphere radius $r_{ps}$ is a physical quantity that continuously depends on the horizon radius $r_{+}$, and its values do not coincide in different branches. In most known black hole solutions, this dependence is in fact monotonic, which yields a one-to-one correspondence between the branches. Consequently, by taking $\mathcal{F}=r_{ps}$, the multivaluedness of $T(r_{+})$ directly implies that of $T(r_{ps})$, and scaling these physical quantities does not change this fundamental property.
@@
As a direct application of this framework, we propose the following geometric criterion for diagnosing a phase transition: a black hole undergoes a first-order phase transition if and only if the temperature curve $T(r_{+})$ exhibits two local extrema. This criterion is equivalent to testing whether $\partial T/\partial r_{+}=0$ has two distinct positive real roots, providing a universal and intuitive tool. Furthermore, we propose a classification scheme that classifies black holes as $A1$, $A2$, and $B$ according to the number of non-degenerate critical points. This classification, based on local geometric properties, forms a complementary perspective to the scheme presented in Ref. Wei et al. (2024), which relies on global topological invariants.
@@ Appendix A Physical nature and mathematical conditions
The mathematical conditions we introduce are grounded in basic physical facts of a first-order phase transition. First, the temperature as a function $T(r_{+})$ is sufficiently smooth, i.e., $T\in C^{p}(I)$ with $p\geq 2$ and $I\subset\mathbb{R}$. More importantly, during a first-order phase transition, this function exhibits two non-degenerate critical points, namely $T^{\prime}(r_{v})=0$ and $T^{\prime\prime}(r_{v})\neq 0$ for $v=1,2$, which correspond to the two endpoints $T_{1}$ and $T_{2}$ of the spinodal region. This is because, when a first-order phase transition occurs, for Maxwell’s equal area law to be meaningful, the function must have two local extremum points with non-zero second derivatives. Hence, for a single-variable function, such extremum points are precisely non-degenerate critical points. Outside the spinodal region, $T(r_{+})$ is strictly monotonic. Finally, $T(r_{+})$ is not a constant function on its entire domain. For a black hole that satisfies the Einstein equations, its Hawking temperature $T(r_{+})$ may inherently satisfy or violate these conditions, which directly determines its phase transition behavior.